### v1

@@ Abstract
This paper presents and investigates an inexact proximal gradient method for solving composite convex optimization problems characterized by an objective function composed of a sum of a full domain differentiable convex function and a non-differentiable convex function. We introduce an explicit linesearch strategy that requires only a relative inexact solution of the proximal subproblem per iteration. We prove the convergence of the sequence generated by our scheme and establish its iteration complexity, considering both the functional values and a residual associated with first-order stationary solutions. Additionally, we provide numerical experiments to illustrate the practical efficacy of our method.
@@ 1 Introduction
Proximal gradient methods effectively solve optimization problems such as in (1). The main step in the proximal gradient method involves evaluating the proximal operator ${\bf prox}_{g}:\mathbbm{E}\to{\bf dom\,}g:=\{x\in\mathbbm{E}\mid g(x)<+\infty\}$ defined by:
${\bf prox}_{g}(y):=\underset{x\in\mathbbm{E}}{\arg\min}\;\left\{g(x)+\frac{1}{2}\left\|x-y\right\|^{2}\right\},$ (2)
@@
Proximal Gradient Method (PGM) Let $x_{0}\in{\bf dom\,}g$. Compute $\gamma_{k}>0$ and $\bar{x}_{k}:={\bf prox}_{\gamma_{k}g}\left(x_{k}-\gamma_{k}\nabla f(x_{k})\right).$ (3) Choose $\beta_{k}\in(0,1]$ and compute $x_{k+1}:=x_{k}+\beta_{k}(\bar{x}_{k}-x_{k}).$ (4)
The coefficients $\gamma_{k}$ and $\beta_{k}$, referred to as stepsizes, can be determined based on backtracking linesearch procedures. Such strategies are essential whenever the global $L$-Lipschitz continuity of $f$ fails or even when computing an acceptable upper bound for $L$ is challenging. This situation is often encountered in numerous applications; for instance, in inverse problems based on Banach norms Bredies:2009 ; Schuster:2012 or Bregman distances such as the Kullback-Leibler divergence Salzo:2014 ; Bonettini:2012 ; BelloCruz:2016a . Moreover, even when $L$ is known, linesearches may allow for longer steps toward the solution by using local information at every iteration.
There are several possible choices for these stepsizes, each impacting the algorithm’s performance in different ways; see, for instance, Salzo:2017a ; BelloCruz:2016a ; BelloCruz:2016c ; Bello-Cruz:2021 . It is important to note that in order to compute the stepsize $\gamma_{k}$ using a backtracking linesearch at each iteration $k$, the proximal operator may need to be evaluated multiple times within the procedure. Conversely, the stepsize $\beta_{k}$ can be selected by evaluating the proximal operator only once per iteration. In this context, we will refer to explicit linesearch to describe a backtracking procedure that determines $\beta_{k}$ after setting $\gamma_{k}$ as a constant for all $k$. These explicit strategies are particularly advantageous, especially in cases where evaluating the proximal operator is challenging. The function $g$ is often complex enough that the corresponding proximal operator lacks an analytical solution. In such cases, an ad-hoc algorithm should be employed to evaluate the proximal operator inexactly. For instance, consider $\mathbb{E}=\mathbbm{R}^{n}$, and let $g:\mathbbm{R}^{n}\to\mathbbm{R}$ be defined as
@@
which will be considered in the numerical illustrations of this paper. For further examples and discussions, see BelloCruz:2017a ; Bello-Cruz:2023 ; Salzo:2012 ; Villa:2013 ; Schmidt:2011 ; Jiang:2012 ; Millan:2019 ; Barre:2022 . Of course, there are rare cases when the exact analytical solution of the proximal operator is available, such as when $g:\mathbbm{R}^{n}\to\mathbbm{R}$ as $g(x)=\|x\|_{1}$ or the indicator function of a simple convex and closed set, see, for instance, Beck:2009 ; Combettes:2005 .
@@
Consequently, in practice, the evaluation of the proximal operator is often done inexactly. A basic inexactness criterion is to approximately evaluate the proximal operator using an exogenous sequence of error tolerance which, in general, must be summable in order to guarantee the convergence of the proximal iterations. Such a diminishing sequence is a priori chosen without using any information that may be available along with the iterations. Usually, the restrictive summability condition forces the solution of the proximal subproblem to be increasingly accurate, often more than necessary; see, for instance, Jiang:2012 ; Salzo:2012 ; Aujol:2015 ; Villa:2013 ; Schmidt:2011 . In the past two decades, relative error criteria have been considered as an effective and practical way of controlling the inexactness in solving proximal subproblems of several algorithms, including FISTA, ADMM, augmented Lagrangian, and Douglas-Rachford methods; refer to Eckstein:2013 ; Eckstein:2018 ; Alves:2020 ; Adona:2019 ; Adona:2020 for examples. Relative error criteria are often easy to verify in practice and has the advantage of exploring the available information at a specific iteration, being, therefore, an interesting alternative to the aforementioned exogenous sequences.
@@
In this paper, we propose and analyze an inexact proximal gradient method for solving problem (1). We propose a novel relative inexactness criterion for solving the proximal subproblems which in some sense ressembles the ideas of relative error criteria introduced in Monteiro:2010 ; Solodov:1999a , but adds some new elements in order to control the objective function on the inexact solution. The proposed scheme requires only one inexact solution of the proximal subproblem per iteration, and the stepsizes are computed through a relaxed explicit linesearch procedure that takes into account the residuals obtained in the proximal subproblem and enables the iteration to address non-Lipschitz optimization problems effectively. We show that the sequence generated by our method converges to a solution of problem (1). Moreover, we establish its iteration-complexity in terms of both the function values and the residues associated to an approximate stationary solution. We also present some numerical experiments to illustrate the performance of the proposed method.
@@
The paper is structured as follows: Section 2 presents definitions, basic facts, and auxiliary results. In Section 3, we introduce the inexact proximal gradient method with an explicit linesearch (IPG-ELS) and establish some of its fundamental properties. Section 4 establishes the iteration-complexity bounds of the IPG-ELS method in terms of functional values and a residual associated with the stationary condition for problem (1). Additionally, the full convergence of the sequence generated by the IPG-ELS method is discussed in this section. Some numerical experiments illustrating the behavior of the proposed scheme are reported in Section 5. Finally, concluding remarks are provided in Section 6.
@@ 2 Preliminary
We now introduce a concept of approximate stationary solution to problem (1). First, note that $\bar{x}$ is a solution to problem (1) if and only if $0\in\nabla f(\bar{x})+\partial g(\bar{x})$. The concept of approximate stationary solution relaxes this inclusion by introducing a residual $v$ and enlarging $\partial g$ using $\partial_{\varepsilon}g$.
@@
Next, we recall the definition of quasi-Fejér convergence, which is an important and well-known tool for establishing full convergence of gradient and proximal point type methods.
@@
The following result presents the main properties of quasi-Fejér convergent sequences; see Theorem $2.6$ of Bauschke:1996 .
@@
If a cluster point $\bar{x}$ of $(x_{k})_{k\in\mathbbm{N}}$ belongs to $S$, then $(x_{k})_{k\in\mathbbm{N}}$ is convergent to $\bar{x}$.

### v2

@@ Abstract
This paper presents and investigates an inexact proximal gradient method for solving composite convex optimization problems characterized by an objective function composed of a sum of a full-domain differentiable convex function and a non-differentiable convex function. We introduce an explicit line search applied specifically to the differentiable component of the objective function, requiring only a relative inexact solution of the proximal subproblem per iteration. We prove the convergence of the sequence generated by our scheme and establish its iteration complexity, considering both the functional values and a residual associated with first-order stationary solutions. Additionally, we provide numerical experiments to illustrate the practical efficacy of our method.
@@ 1 Introduction
Proximal gradient methods effectively solve optimization problems such as (1). The main step in the proximal gradient method involves evaluating the proximal operator ${\bf prox}_{g}:\mathbbm{E}\to{\bf dom\,}g:=\{x\in\mathbbm{E}\mid g(x)<+\infty\}$ defined as follows:
${\bf prox}_{g}(x):=\underset{y\in\mathbbm{E}}{\arg\min}\;\left\{g(y)+\frac{1}{2}\left\|y-x\right\|^{2}\right\},$ (2)
@@
Proximal Gradient Method (PGM) Let $x_{0}\in{\bf dom\,}g$. Compute $\lambda_{k}>0$ and $\tilde{x}_{k}:={\bf prox}_{\lambda_{k}g}\left(x_{k}-\lambda_{k}\nabla f(x_{k})\right).$ (3) Choose $\beta_{k}\in(0,1]$ and compute $x_{k+1}:=x_{k}+\beta_{k}(\tilde{x}_{k}-x_{k}).$ (4)
The coefficients $\lambda_{k}$ and $\beta_{k}$, referred to as stepsizes, can be determined based on backtracking linesearch procedures. Such strategies are essential whenever the global $L$-Lipschitz continuity of the gradient of $f$ fails or even when computing an acceptable upper bound for $L$ is challenging. This situation is often encountered in numerous applications; for instance, in inverse problems based on non-Euclidean norms Bredies:2009 ; Schuster:2012 or Bregman distances such as the Kullback-Leibler divergence Salzo:2014 ; Bonettini:2012 ; Csiszar:1991 ; Vardi:1985 . Moreover, even when $L$ is known, linesearches may allow for longer steps toward the solution by using local information at every iteration.
There are several possible choices for these stepsizes, each impacting the algorithm’s performance in different ways; see, for instance, Salzo:2017a ; BelloCruz:2016a ; BelloCruz:2016c ; Bello-Cruz:2021 . It is important to note that in order to compute the stepsize $\lambda_{k}$ using a backtracking linesearch at each iteration $k$, the proximal operator may need to be evaluated multiple times within the procedure. Conversely, the stepsize $\beta_{k}$ can be selected by evaluating the proximal operator only once per iteration. In this context, we will refer to explicit linesearch to describe a backtracking procedure that determines $\beta_{k}$ after setting $\lambda_{k}$ as a constant for all $k$. This type of explicit strategy, first presented in BelloCruz:2016a , is particularly advantageous, especially in cases where evaluating the proximal operator is challenging. The function $g$ is often complex enough that the corresponding proximal operator lacks an analytical solution. In such cases, an ad-hoc algorithm should be employed to evaluate the proximal operator inexactly. For instance, consider $\mathbb{E}=\mathbbm{R}^{n}$, and let $g:\mathbbm{R}^{n}\to\mathbbm{R}$ be defined as
@@
where $\lambda_{\text{row}},\lambda_{\text{col}}>0$. This nondifferentiable term in problem (1) will be considered in the numerical illustrations of this paper. For further examples and discussions, see BelloCruz:2017a ; Bello-Cruz:2023 ; Salzo:2012 ; Villa:2013 ; Schmidt:2011 ; Jiang:2012 ; Millan:2019 ; Barre:2022 ; BelloCruz:2022b . Of course, there are some cases when the exact analytical solution of the proximal operator is available, such as when $g:\mathbbm{R}^{n}\to\mathbbm{R}$ as $g(x)=\|x\|_{1}$ or the indicator function of a simple convex and closed set, see, for instance, Beck:2009 ; Combettes:2005 .
@@
Consequently, in practice, the evaluation of the proximal operator is often done inexactly. A basic inexactness criterion is to approximately evaluate the proximal operator using an exogenous sequence of error tolerances which, in general, must be summable in order to guarantee the convergence of the proximal iterations. Such a diminishing sequence is a priori chosen without using any information that may be available along with the iterations. Usually, the restrictive summability condition forces the solution of the proximal subproblem to be increasingly accurate, often more than necessary; see, for instance, Jiang:2012 ; Salzo:2012 ; Aujol:2015 ; Villa:2013 ; Schmidt:2011 . In the past two decades, relative error criteria have been considered an effective and practical way of controlling the inexactness in solving proximal subproblems of several algorithms, including FISTA, ADMM, augmented Lagrangian, Douglas-Rachford, and proximal gradient methods; refer to Eckstein:2013 ; Eckstein:2018 ; Alves:2020 ; Adona:2019 ; Adona:2020 ; Bonettini:2016 ; Bello-Cruz:2023 ; Millan:2019 for examples. Relative error criteria are often easy to verify in practice and have the advantage of exploiting the available information at a specific iteration, being, therefore, an interesting alternative to the aforementioned exogenous sequences.
@@
In this paper, we propose and analyze an inexact proximal gradient method (PGM) for solving problem (1). We present a novel relative inexactness criterion for solving the proximal subproblems which somewhat resembles the ideas of relative error criteria introduced in Monteiro:2010 ; Solodov:1999a , but incorporates some new elements to control the objective function for the inexact solution. The proposed scheme requires only one inexact solution of the proximal subproblem per iteration, and the stepsizes are computed through a relaxed explicit linesearch procedure, applied specifically to $f$, that takes into account the residuals obtained from the proximal subproblem and enables the iteration to address non-Lipschitz optimization problems effectively. We show that the sequence generated by our method converges to a solution of problem (1). Moreover, we establish its iteration complexity in terms of both the function values and the residuals associated with an approximate stationary solution. We also present some numerical experiments to illustrate the performance of the proposed method.
It is worth mentioning that the PGM proposed in this paper can be regarded as an inexact version of the ones analyzed in Salzo:2017a ; BelloCruz:2016a , where explicit linesearches were studied in the PGM setting. On the other hand, an inexact version of the PGM with an explicit linesearch was also proposed in Bonettini:2016 . The authors developed a general framework based on variable metrics and established results on convergence rates for the sequence generated by their method. While, in the special case where the metric is fixed and chosen as the Euclidean one, their algorithm shares some similarities with ours, it employs different type of linesearch and inexact criteria for solving the proximal subproblem. Further details on the differences between our algorithm and the method in Bonettini:2016 are discussed in Subsection 3.3.
@@
The paper is structured as follows: Section 2 presents definitions, basic facts, and auxiliary results. The concept of an approximate solution for the proximal subproblem, along with the description of algorithms that can be employed to compute it, is detailed in Section 3. In this latter section, we discuss some works that analyze relative inexact proximal solution criteria related to ours. Section 4 introduces the inexact proximal gradient method with an explicit linesearch (IPG-ELS) and establishes some of its fundamental properties. Section 5 analyzes the full convergence of the sequence generated by the IPG-ELS method and establishes its iteration complexity bounds in terms of functional values and a residual associated with the stationary condition for problem (1). Some numerical experiments illustrating the behavior of the proposed scheme are reported in Section 6. Finally, concluding remarks are provided in Section 7.
@@ 2 Preliminary
We now introduce a concept of approximate stationary solution to problem (1), which can be seen as a specialization of the one presented in (Monteiro:2010, , Eq. 1). First, note that $\bar{x}$ is a solution to problem (1) if and only if $0\in\nabla f(\bar{x})+\partial g(\bar{x})$. The concept of approximate stationary solution relaxes this inclusion by introducing a residual $v$ and enlarging $\partial g$ using $\partial_{\varepsilon}g$.
@@
Next, we recall the definition of quasi-Fejér convergence, which is an important and well-known tool for establishing full convergence of gradient and proximal point type methods; see, for instance, (Iusem-Quasi-Fejer, , Definition 1).
@@
The following result presents the main properties of quasi-Fejér convergent sequences; see (Bauschke:1996, , Theorem $2.6$).
@@
If an accumulation point $\bar{x}$ of $(x_{k})_{k\in\mathbbm{N}}$ belongs to $S$, then $(x_{k})_{k\in\mathbbm{N}}$ is convergent to $\bar{x}$.